4. Two identical objects move with speeds of 5.0 m/s and 25.0 m/s. What is the ratio of their kinetic energies?
Answer #1

As mentioned in the problem, objects are identical. So, their mass will be equal.

Suppose mass of the each object = m kg.

Expression for the kinetic energy of an object of mass ‘m’ and velocity ‘v’ is given as –

KE = (1/2)*m*v^2

For first mass, v1 = 5.0 m/s

So, Kinetic energy, KE1 = (1/2)*m*v1^2 = (1/2)*m*5.0^2 Joule.

For second mass, v2 = 25.0 m/s

So, Kinetic energy, KE2 = (1/2)*m*v2^2 = (1/2)*m*25.0^2 Joule.

Therefore ratio of their kinetic energies –

KE1 / KE2 = [(1/2)*m*5.0^2] / [(1/2)*m*25.0^2] = 25.0 / 625.0 = 1/25.0 = 0.04 (Answer)